Grade 12Mechanics

16 gm of o2 expand at constant pressure and 273.15K to occupy double it's original volume . The work done during the process will be

JHS
6 Years agoGrade 12

1 Answer

Arun
6 Years ago
 
Since final volume is twice that of initial volume so 
W= P*(Vfinal-Vinitial) 
W=1*(2V-V)
W=V
Now no. Of moles = 16000/32= 500 
PV=nRT 
V=500*2*273=27300
Now W =27300/1000= 273kcal